# Unleashing Diffusion Transformers for Visual Correspondence by Modulating Massive Activations

### v1

@@ Abstract
Pre-trained stable diffusion models (SD) have shown great advances in visual correspondence. In this paper, we investigate the capabilities of Diffusion Transformers (DiTs) for accurate dense correspondence. Distinct from SD, DiTs exhibit a critical phenomenon in which very few feature activations exhibit significantly larger values than others, known as massive activations, leading to uninformative representations and significant performance degradation for DiTs. The massive activations consistently concentrate at very few fixed dimensions across all image patch tokens, holding little local information. We trace these dimension-concentrated massive activations and find that such concentration can be effectively localized by the zero-initialized Adaptive Layer Norm (AdaLN-zero). Building on these findings, we propose Diffusion Transformer Feature (DiTF), a training-free framework designed to extract semantic-discriminative features from DiTs. Specifically, DiTF employs AdaLN to adaptively localize and normalize massive activations with channel-wise modulation. In addition, we develop a channel discard strategy to further eliminate the negative impacts from massive activations. Experimental results demonstrate that our DiTF outperforms both DINO and SD-based models and establishes a new state-of-the-art performance for DiTs in different visual correspondence tasks (e.g., with +9.4% on Spair-71k and +4.4% on AP-10K-C.S.).
@@ 1 Introduction
Detecting dense correspondence between images plays an important role in various computer vision tasks such as segmentation liu2021unsupervised ; hamilton2022unsupervised ; liu2023matcher ; khani2023slime , image editing alaluf2024cross ; zhang2020cross ; go2024eye , and point tracking yan2022towards ; karaev2023cotracker ; wang2023tracking ; tumanyan2025dino . Various works oquab2023dinov2 ; darcet2023vision ; zhang2024tale utilize pre-trained self-supervised models like DINOv2 oquab2023dinov2 to extract image features, which are then adapted for precise semantic correspondence. With the advent of high fidelity text-to-image (T2I) generative models rombach2022high ; saharia2022photorealistic , recent studies hedlin2024unsupervised ; tang2023emergent ; zhang2024tale have demonstrated the powerful capability of pre-trained SD rombach2022high as an effective feature extractor zhang2024tale ; tang2023emergent ; li2024sd4match for accurate visual correspondence.
@@
However, when directly extracting the original features for representation learning, as done in SD, DiTs exhibit disappointing performance for visual perception tasks, as shown in Figure 1. To better understand this discrepancy, we analyze the statistical properties of features extracted from both SD and DiTs (see Figure 2). Notably, we observe that DiT feature maps exhibit massive activations, in contrast to the smoother distribution observed in SD2.1 rombach2022high . These massive activations share a few fixed dominant dimensions, resulting in a high degree of directional similarity for feature vectors in the latent space lan2024clearclip . This property makes it difficult to distinguish spatial feature vectors using cosine similarity, leading to semantically uninformative and indiscriminate representations.
In this work, we aim to better understand the nature of massive activations in DiTs and extract semantically discriminative representations from DiTs for visual correspondence tasks. We begin by investigating the spatial and dimensional properties of these activations and find that they consistently emerge in a small set of fixed feature dimensions across all image patch tokens, which carry limited local information that are crucial for our task. Further analysis reveals that these concentrated dimensions align with the residual scaling factors generated by the AdaLN-zero layers. Motivated by this observation, we examine the role of AdaLN-zero in modulating DiT features. Our findings reveal that the built-in AdaLN-zero effectively localizes massive activations and performs semantically meaningful channel-wise modulation to suppress them, thereby enhancing the semantic richness and discriminative power of DiT features.
Based on these findings, we propose a training-free framework: Diffusion Transformer Feature (DiTF), to extract structural and semantic features from DiTs. Specifically, DiTF adaptively normalizes massive activations with channel-wise modulation by the built-in AdaLN in the DiT block. In addition, we develop a channel discard strategy to further eliminate the undesired impacts from massive activations for representation learning. Our main contributions can be summarized as follows.
  • We are the first to identify and characterize massive activations in Diffusion Transformers (DiTs). These activations are consistently concentrated in a small number of fixed feature dimensions across all image patch tokens and encode limited local information.
  • We trace the dimension-concentrated massive activations and find that such concentration can be effectively localized by the built-in AdaLN-zero layer in DiTs.
  • We propose a training-free framework, Diffusion Transformer Feature (DiTF), for extracting semantically discriminative representations from DiTs.
@@ 2 Related Work
Massive Activations. Massive activations have been extensively studied in large language models (LLMs) sun2024massive ; zhao2023unveiling ; xu2025slmrec . Several works sun2024massive ; xiao2023efficient have shown that these activations are concentrated in a small number of fixed dimensions, particularly in the start and delimiter tokens. jin2025massive further revealed that the massive values in the query and key vectors originate from the positional encoding mechanism RoPEsu2024roformer . In addition to LLMs, similar phenomena have also been observed in Vision Transformers (ViTs)darcet2023vision ; yang2024denoising ; sun2024massive . Specifically, sun2024massive identified attention artifacts in ViT feature maps, predominantly appearing in low-information background regions. DVT yang2024denoising further traced these artifacts to the influence of positional embeddings.
@@ 3 Preliminaries
where $z_{t}^{(k,1)}$ and $z_{t}^{(k+1,1)}$ denote the same features as $z_{t}^{k}$ and $z_{t}^{k+1}$, respectively. For clarity, we refer to $z_{t}^{(k,1)}$ and $z_{t}^{(k,2)}$ as the pre-AdaLN intermediate features, and $\hat{z}_{t}^{(k,1)}$ and $\hat{z}_{t}^{(k,2)}$ as the post-AdaLN intermediate features in the subsequent sections. The residual scaling factor $\alpha_{k}$ in Equation 2 corresponds to $\alpha_{k}^{2}$. For simplicity, we use AdaLN to refer to the AdaLN-zero module in DiTs throughout this paper. See the Appendix A for an illustration of DiTs architecture peebles2023scalable .
@@ 4.1 Massive Activations in Diffusion Transformers
We identify the dimensional locations of massive activations within the original DiT feature. As shown in Figure 3, the SD3-5 model exhibits massive activations concentrated in a single feature dimension (676). Similarly, in the Flux model, massive activations appear in two specific dimensions (154 and 1446). Analyses of other DiTs are provided in the Appendix. These observations suggest that massive activations consistently occur in a small number of fixed feature dimensions, which aligns with similar findings in LLMs sun2024massive .
Massive activations appear across all patch tokens.
As shown in Figures 2 and 3, massive activations consistently appear across all image patch tokens, regardless of the input image. For instance, in the SD3-5 model, the 676th feature dimension exhibits massive activations across all patch tokens. This phenomenon differs from the observations in LLMs and Vision Transformers (ViTs). Specifically, sun2024massive reported that massive activations in LLMs are primarily concentrated in special tokens such as start and delimiter tokens, while darcet2023vision found that outlier activations in ViTs tend to emerge in low-information tokens.
@@
To better understand the nature of massive activations, we analyze their spatial variance by measuring the mean and variance of feature values across image patches. Feature dimensions are ranked by their averaged activation magnitude across patches. We then compare the top-ranked dimension (massive activations) with the 10th and 20th highest-ranked (non-massive) dimensions to assess their spatial discriminability. As shown in Table 1, the variance of the massive activations (1st) is substantially lower relative to its mean compared to the non-massive activations (10th and 20th), indicating that massive activations are less spatially discriminative. This suggests that massive activations encode limited local information.
@@
It can be seen that the original features in DiTs are computed as $z_{t}^{k+1}=z_{t}^{k}+\alpha_{k}\mathcal{A}_{k}(z_{t}^{k})$ via residual connections, where a scaling factor $\alpha_{k}$ is applied to weight the block output $\mathcal{A}_{k}(z_{t}^{k})$ (see Equation 2). To examine the potential link between massive activations and the residual scaling factor $\alpha_{k}$, we analyze their dimensional correspondence in Figure 4 and observe that they consistently co-occur in the same channels. Notably, $\alpha_{k}$ is regressed by the AdaLN-zero layer, suggesting that AdaLN-zero enables accurate localization of massive activations.
@@ 4.2 Channel-wise Modulation with AdaLN
To better understand how AdaLN interacts with massive activations, we compare the properties of pre-AdaLN and post-AdaLN intermediate features within DiT blocks. As shown in Figure 5, massive activations in feature channels are significantly reduced after AdaLN processing. This suggests that AdaLN can effectively identify such concentrated activations and adaptively normalize them through channel-wise modulation, using the scaling and shifting parameters.
@@
Furthermore, we visualize different feature maps in Figure 5. It can be observed that the pre-AdaLN feature maps in DiTs exhibit limited semantic coherence, primarily capturing low-level textures such as object color (e.g., the cow and bird), and show weak discrimination between objects and background. In contrast, the post-AdaLN features display markedly improved spatial coherence and clearer semantic boundaries across object parts, suggesting that AdaLN improves spatial semantics and feature discrimination. More visualization results can be found in the Appendix, where the same phenomenon is observed.
@@ 4.3 Extracting features from Diffusion Transformers
As shown in Figure 5, it can be observed that the post-AdaLN features still exhibit a few weakly massive activations, which will slightly compromise the local semantics of feature maps. Therefore, we propose a simple yet effective channel discard strategy to further suppress their influence. Specifically, we zero out the corresponding dimensions of these weakly massive activations in the post-AdaLN features as they hold little local information.
@@
[…]
@@ 5.2 Main results on semantic correspondence
[…]
Cross-category Semantic Correspondence In addition, we conducted cross-category semantic correspondence experiments on the AP-10K dataset, as shown in Table 3. The results indicate that our models enable the extraction of accurate semantic features for cross-category image pairs and the drawing of correct correspondences, outperforming the previous best method, GeoAware-SC zhang2024telling , by nearly 4.8% (64.6% to 69.4%)on AP-10K-C.S. Furthermore, the results show that previous methods, such as DINOv2 and DIFT, experience significant performance degradation on the cross-category task. In contrast, our model demonstrates stronger robustness for the cross-category correspondence.
@@
[…]
@@ 5.3 Further analysis
Evaluation of Channel Modulation with AdaLN. We conducted comprehensive semantic correspondence experiments across different datasets on various DiTs. The results, as depicted in Figure 6 and Table 4, reveal several key observations: 1) Effectiveness. When integrated with the channel modulation by AdaLN, all the DiT models experience an absolute performance boost of more than 20%, demonstrating that AdaLN is capable of effectively mitigating massive activations and enhancing feature semantics and discrimination.  2) Robustness. These results indicate that channel modulation with AdaLN enables the accurate normalization of features across different categories, leading to significant improvements on cross-category datasets.
Evaluation of Channel Discard strategy. As shown in Figure 6 and Table 4. We can observe that our channel discard strategy can effectively locate and eliminate the weakly massive activations in the DiT feature across different DiTs, leading to smoother feature maps and higher correspondence matching accuracy.
Impact of conditions in AdaLN for Channel-wise Modulation. From the results in Table 5, we can observe that the timestep $t$ is essential for modulating massive activations, bringing a significant performance boost. In contrast, attempting to modulate features solely based on $c$ proves ineffective. This phenomenon arises from the inherent nature of DiTs. During training, DiTs introduce noise to images at varying levels of $t$, making the timestep an essential condition for feature modulation. It serves as a key factor in locating the channel positions of massive activations and enhancing feature quality, enabling their transformations into a clean and semantically meaningful representation.
Qualitative Results. We visualize some semantic correspondence predictions from different models in Figures 7 and 8. The visualization reveals that our DiT-based model yields robust and accurate correspondences across various complex scenes, including changes in viewpoint, multiple objects, and dense matching. These results demonstrate the superiority and robustness of our models.

### v3

@@ Abstract
Pre-trained stable diffusion models (SD) have shown great advances in visual correspondence. In this paper, we investigate the capabilities of Diffusion Transformers (DiTs) for accurate dense correspondence. Distinct from SD, DiTs exhibit a critical phenomenon in which very few feature activations exhibit significantly larger values than others, known as massive activations, leading to uninformative representations and significant performance degradation for DiTs. The massive activations consistently concentrate at very few fixed dimensions across all image patch tokens, holding little local information. We analyze these dimension-concentrated massive activations and uncover that their concentration is inherently linked to the Adaptive Layer Normalization (AdaLN) in DiTs. Building on these findings, we propose the Diffusion Transformer Feature (DiTF), a training-free AdaLN-based framework that extracts semantically discriminative features from DiTs. Specifically, DiTF leverages AdaLN to adaptively localize and normalize massive activations through channel-wise modulation. Furthermore, a channel discard strategy is introduced to mitigate the adverse effects of massive activations. Experimental results demonstrate that our DiTF outperforms both DINO and SD-based models and establishes a new state-of-the-art performance for DiTs in different visual correspondence tasks (e.g., with +9.4% on Spair-71k and +4.4% on AP-10K-C.S.).
@@ 1 Introduction
Detecting dense correspondence between images plays an important role in various vision tasks such as segmentation liu2021unsupervised; hamilton2022unsupervised; liu2023matcher; khani2023slime, image editing alaluf2024cross; zhang2020cross; go2024eye, and point tracking yan2022towards; karaev2023cotracker; wang2023tracking; tumanyan2025dino. Various works oquab2023dinov2; darcet2023vision; zhang2024tale utilize pre-trained self-supervised models like DINOv2 oquab2023dinov2 to extract image features, which are then adapted for precise semantic correspondence. With the advent of text-to-image (T2I) generative models rombach2022high; saharia2022photorealistic, recent studies hedlin2024unsupervised; tang2023emergent; zhang2024tale have demonstrated the powerful capability of pre-trained SD rombach2022high as an effective feature extractor zhang2024tale; tang2023emergent; li2024sd4match for visual correspondence.
@@
However, when directly extracting the original features for representation learning, as done in SD, DiTs exhibit disappointing performance for visual perception tasks, as shown in Figure˜1. To better understand this discrepancy, we analyze the statistical properties of features extracted from both SD and DiTs (see Figure˜2). Notably, we observe that DiT feature maps exhibit massive activations, in contrast to the smoother distribution observed in SD2.1 rombach2022high. These massive activations share a few fixed dominant dimensions, resulting in a high degree of directional similarity for feature vectors in the latent space lan2024clearclip. This property makes it difficult to distinguish spatial feature vectors using cosine similarity, leading to semantically uninformative and indiscriminate representations.
In this work, we aim to deepen the understanding of massive activations for representation learning, thereby enabling the extraction of semantically discriminative representations from DiTs. We begin by investigating the spatial and dimensional properties of these activations and find that they consistently emerge in a small set of fixed feature dimensions across all image tokens, which carry limited local information that is crucial for our task. Further analysis reveals an intriguing connection between massive activations and the AdaLN in DiTs, where the concentrated dimensions of massive activations align with the residual scaling factors produced by AdaLN. To further understand this phenomenon, we investigate the role of AdaLN in modulating DiT features. Our findings reveal that the built-in AdaLN effectively localizes massive activations and performs semantically meaningful channel-wise modulation to suppress them, thereby enhancing the semantic richness and discriminative capacity of DiT features.
Based on these findings, we propose a training-free AdaLN-based framework: Diffusion Transformer Feature (DiTF), to extract structural and semantic features from DiTs. Specifically, DiTF leverages the built-in AdaLN in DiTs to adaptively normalize massive activations through channel-wise modulation. In addition, a channel discard strategy is introduced to further mitigate the adverse effects of massive activations on representation learning. Our main contributions are summarized as follows.
  • We identify and characterize massive activations in DiTs and find them consistently concentrated in a small number of fixed feature dimensions across all spatial tokens.
  • We analyze the source of these dimensionally concentrated massive activations and uncover their strong connection to the AdaLN layer in DiTs.
  • We propose a training-free AdaLN-based framework, Diffusion Transformer Feature (DiTF), for extracting semantically discriminative representations from DiTs.
@@ 2 Related Work
Massive Activations. Massive activations have been extensively studied in large language models (LLMs) sun2024massive; zhao2023unveiling; xu2025slmrec. Several works sun2024massive; xiao2023efficient have shown that these activations are concentrated in a small number of fixed dimensions, particularly in the start and delimiter tokens. jin2025massive further revealed that the massive values in the query and key vectors originate from the positional encoding mechanism RoPEsu2024roformer. In addition to LLMs, similar phenomena have also been observed in Vision Transformers (ViTs)darcet2023vision; yang2024denoising; sun2024massive. Specifically, sun2024massive identified attention artifacts in ViT feature maps, predominantly appearing in low-information background regions. DVT yang2024denoising further traced these artifacts to the influence of positional embeddings. Recent studies on accelerating DiTs fang2025tinyfusion; zhao2025viditq have identified massive activation outliers in DiTs and revealed that these outliers lead to unstable model quantization and excessively high distillation losses during knowledge distillation.
@@ 3 Preliminaries
where $z_{t}^{(k,1)}$ and $z_{t}^{(k+1,1)}$ denote the same features as $z_{t}^{k}$ and $z_{t}^{k+1}$, respectively. For clarity, we refer to $z_{t}^{(k,1)}$ and $z_{t}^{(k,2)}$ as the pre-AdaLN intermediate features, and $\hat{z}_{t}^{(k,1)}$ and $\hat{z}_{t}^{(k,2)}$ as the post-AdaLN intermediate features in the subsequent sections. The residual scaling factor $\alpha_{k}$ in Equation˜2 corresponds to $\alpha_{k}^{2}$. For simplicity, we use AdaLN to refer to the AdaLN-zero module in DiTs throughout this paper. See the Appendix A for an illustration of DiTs architecture peebles2023scalable.
@@ 4.1 Massive Activations in Diffusion Transformers
We identify the dimensional locations of massive activations within the original DiT feature. As shown in Figure˜3, the SD3-5 model exhibits massive activations concentrated in a single feature dimension (676). Similarly, in the Flux model, massive activations appear in two specific dimensions (154 and 1446). Analyses of other DiTs are provided in the Appendix. These observations suggest that massive activations consistently occur in a small number of fixed feature dimensions, which aligns with similar findings in LLMs sun2024massive.
Massive activations appear across all spatial tokens.
As shown in Figures˜2 and 3, massive activations consistently appear across all image patch tokens, regardless of the input image. For instance, in the SD3-5 model, the 676th feature dimension exhibits massive activations across all patch tokens. This phenomenon differs from the observations in LLMs and Vision Transformers (ViTs). Specifically, sun2024massive reported that massive activations in LLMs are primarily concentrated in special tokens such as start and delimiter tokens, while darcet2023vision found that outlier activations in ViTs tend to emerge in low-information tokens.
@@
To better understand the nature of massive activations, we analyze their spatial variance by measuring the mean and variance of feature values across image patches. Feature dimensions are ranked by their averaged activation magnitude across patches. We then compare the top-ranked dimension (massive activations) with the 10th and 20th highest-ranked (non-massive) dimensions to assess their spatial discriminability. As shown in Table˜1, the variance of the massive activations (1st) is substantially lower relative to its mean compared to the non-massive activations (10th and 20th), indicating that massive activations are less spatially discriminative. This suggests that massive activations encode limited local information.
@@
It can be seen that the original features in DiTs are computed as $z_{t}^{k+1}=z_{t}^{k}+\alpha_{k}\mathcal{A}_{k}(z_{t}^{k})$ via residual connections, where a scaling factor $\alpha_{k}$ is applied to weight the block output $\mathcal{A}_{k}(z_{t}^{k})$ (see Equation˜2). To examine the potential link between massive activations and the residual scaling factor $\alpha_{k}$, we analyze their dimensional correspondence in Figure˜4 and observe that they consistently co-occur in the same channels. Notably, $\alpha_{k}$ is regressed by the AdaLN-zero layer, suggesting that AdaLN-zero enables accurate localization of massive activations.
@@ 4.2 Channel-wise Modulation with AdaLN
To better understand how AdaLN interacts with massive activations, we compare the properties of pre-AdaLN and post-AdaLN intermediate features within DiT blocks. As shown in Figure˜5, massive activations in feature channels are significantly reduced after AdaLN processing. This suggests that AdaLN can effectively identify such concentrated activations and adaptively normalize them through channel-wise modulation, using the scaling and shifting parameters.
@@
Furthermore, we visualize different feature maps in Figure˜5. It can be observed that the pre-AdaLN feature maps in DiTs exhibit limited semantic coherence, primarily capturing low-level textures such as object color (e.g., the cow and bird), and show weak discrimination between objects and background. In contrast, the post-AdaLN features display markedly improved spatial coherence and clearer semantic boundaries across object parts, suggesting that AdaLN improves spatial semantics and feature discrimination. More visualization results can be found in the Appendix, where the same phenomenon is observed.
@@ 4.3 Extracting features from Diffusion Transformers
As shown in Figure˜5, it can be observed that the post-AdaLN features still exhibit a few weakly massive activations, which will slightly compromise the local semantics of feature maps. Therefore, we propose a simple yet effective channel discard strategy to further suppress their influence. Specifically, we zero out the corresponding dimensions of these weakly massive activations in the post-AdaLN features as they hold little local information.
@@
$\text{DiTF}_{\texttt{sd3-5}}\text{(ours)}$} $\underline{7.9}$ $\underline{48.1}$ $\mathbf{61.2}$ $\mathbf{7.6}$ $\mathbf{50.2}$ $\mathbf{63.3}$ $\underline{6.6}$ $\underline{48.0}$ $\underline{60.2}$ $\mathbf{5.8}$ $\underline{33.5}$ $\underline{47.9}$ $\underline{88.5}$ $\underline{95.2}$ $\underline{97.3}$
$\text{DiTF}_{\texttt{flux}}\text{(ours)}$} $\mathbf{8.3}$ $\mathbf{49.3}$ $\mathbf{64.0}$ $\mathbf{7.6}$ $\underline{49.7}$ $\underline{62.2}$ $\mathbf{6.8}$ $\mathbf{48.2}$ $\mathbf{61.7}$ $\underline{5.7}$ $\mathbf{34.2}$ $\mathbf{48.9}$ $\mathbf{89.5}$ $\mathbf{95.8}$ $\mathbf{97.6}$
@@ 5.2 Main results on semantic correspondence
In-category Semantic Correspondence We conducted comprehensive experiments across different datasets where the results can be found in Tables˜2 and 3. From these results, we could obtain the following observations: 1) State-of-the-art performance. Overall, our models achieve state-of-the-art performance across different datasets. Specifically, our model $\text{DiTF}_{sd3-5}\text{(ours)}$$^{\dagger}$ achieves 2.6% $\uparrow$ on Spair-71k and 2.6% $\uparrow$ on AP-10K-I.S. 2) Superior feature extraction capabilities. It is readily apparent that our feature extraction models for DiTs, such as $\text{DiTF}_{\texttt{sd3-5}}\text{(ours)}$} and $\text{DiTF}_{\texttt{flux}}\text{(ours)}$}, outperform traditional feature extractors like SD2-1 (DIFT tang2023emergent) and DINOv2 oquab2023dinov2 across various datasets, demonstrating their strong and robust feature extraction abilities. Specifically, our $\text{DiTF}_{\texttt{flux}}\text{(ours)}$} achieves a performance of 67.1% (9.4% $\uparrow$ over DIFT) on the Spair-71k dataset, and obtains a performance of 62.2% (1.3% $\uparrow$ over DINOv2) on dataset AP-10K-I.S. 3) Feature redundancy across DiTs and DINOv2: As shown in Table˜2, when integrating DINOv2, our $\text{DiTF}_{\texttt{sd3-5}}$ achieves a notable performance increase (from 64.6% to 72.2%). In comparison, our $\text{DiTF}_{\texttt{flux}}$ shows a limited improvement, only a 0.5% gain (from 67.1% to 67.6%). We interpret this performance discrepancy from feature alignment. FLUX aligns closely with DINOv2, causing redundancy, while SD3.5 benefits more from DINOv2’s complementary distribution. This alignment may arise from two factors: Improved alignment between diffusion models and DINOv2 with larger models and longer training yu2024representation; Overlap in training data.
Cross-category Semantic Correspondence In addition, we conducted cross-category semantic correspondence experiments on the AP-10K dataset, as shown in Table˜3. The results indicate that our models enable the extraction of accurate semantic features for cross-category image pairs and the drawing of correct correspondences, outperforming the previous best method, GeoAware-SC zhang2024telling, by nearly 4.8% (64.6% to 69.4%)on AP-10K-C.S. Furthermore, the results show that previous methods, such as DINOv2 and DIFT, experience significant performance degradation on the cross-category task. In contrast, our model demonstrates stronger robustness for the cross-category correspondence.
@@
$\text{DiTF}_{\texttt{flux}}\text{(w/o AdaLN)}$ 43.8 44.8
@@ 5.3 Further analysis
Evaluation of Channel Modulation with AdaLN. We conducted comprehensive semantic correspondence experiments across different datasets on various DiTs. The results, as depicted in Figure˜6 and Table˜4, reveal several key observations: 1) Effectiveness. When integrated with the channel modulation by AdaLN, all the DiT models experience an absolute performance boost of more than 20%, demonstrating that AdaLN is capable of effectively mitigating massive activations and enhancing feature semantics and discrimination.  2) Robustness. These results indicate that channel modulation with AdaLN enables the accurate normalization of features across different categories, leading to significant improvements on cross-category datasets.
Evaluation of Channel Discard strategy. As shown in Figure˜6 and Table˜4. We can observe that our channel discard strategy can effectively locate and eliminate the weakly massive activations in the DiT feature across different DiTs, leading to smoother feature maps and higher correspondence matching accuracy.
Impact of conditions in AdaLN for Channel-wise Modulation. From the results in Table˜5, we can observe that the timestep $t$ is essential for modulating massive activations, bringing a significant performance boost. In contrast, attempting to modulate features solely based on $c$ proves ineffective. This phenomenon arises from the inherent nature of DiTs. During training, DiTs introduce noise to images at varying levels of $t$, making the timestep an essential condition for feature modulation. It serves as a key factor in locating the channel positions of massive activations and enhancing feature quality, enabling their transformations into a clean and semantically meaningful representation.
Qualitative Results. We visualize some semantic correspondence predictions from different models in Figures˜7 and 8. The visualization reveals that our DiT-based model yields robust and accurate correspondences across various complex scenes, including changes in viewpoint, multiple objects, and dense matching. These results demonstrate the superiority and robustness of our models.